A machine with velocity ratio of 5 was used to rise a load of 1.2kg mass through a hight of 10m, if the work done was 800j. finde the efficiency of the machine
1
Expert's answer
2019-12-18T11:28:10-0500

The mechanical advantage of the machine


"MA=\frac{W}{mgh}=\frac{800}{(1.2)(9.8)(10)}=6.8"

The efficiency of the machine:


"\eta=\frac{MA}{VR}=\frac{6.8}{5}=1.36\ or \ 136\%"